OpenStudy (anonymous):

Solve 3x – 4y = –12 for y.

OpenStudy (anonymous):

algebra 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you have to subtract 3x from both sides 3x-4y=-12 -3x -3x -4y=-3x-12 now divide both sides by -4 -4y=-3x-12 /-4 /-4 y= 3/4x+3

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you welcome